OpenStudy (anonymous):

how do I find the domain of the square root of ( x^2 - 3)

OpenStudy (anonymous):

You can't take the square root of a negative number and get a real result, so solve: \[x^2 - 3 \ge 0\]

OpenStudy (calculusfunctions):

Given an equation, the domain is the set of all values of x for which the equation is defined in the reals. \[y =\sqrt{x ^{2}-3}\] This function has a square root. We know that a square root of a negative number is not defined over reals. Thus the radicand (the expression under the root) must be greater than or equal to zero. Thus x² − 3 ≥ 0 x² ≥ 3 \[x \le -\sqrt{3}, x \ge \sqrt{3}\] And this is your domain. Does that help?

OpenStudy (anonymous):

Yes I does. Thank you both for helping me.